OpenStudy (anonymous):

How to solve for x, y, and z by eliminating: x=1-y 2x=z 2z=-2-y

OpenStudy (anonymous):

Okay. make x the subject for the second equation for me please.

OpenStudy (anonymous):

Wait you want to do this by elimination?

OpenStudy (anonymous):

yes, eliminate the equations until I have solved for one x, y, and

OpenStudy (anonymous):

Yeah okay.

OpenStudy (anonymous):

Move the -y in the first equation to the other side.

OpenStudy (anonymous):

Do the same thing in the third equaiton. Move the -y to the other side.

OpenStudy (anonymous):

Always keep the variables on one side of the equation.

OpenStudy (anonymous):

variables are the pronumerals or letters: x, y and z.

OpenStudy (anonymous):

Give me your rewritten equations for me please after you've done that step.

OpenStudy (anonymous):

x+y=1 2x=z 2z+y=-2

OpenStudy (anonymous):

What about the second equation? Keep ALL variables to one side.

OpenStudy (anonymous):

2x-z=0?

OpenStudy (anonymous):

Yep Good work.

OpenStudy (anonymous):

Okay so now we have these three equations: \[x+y=1\] \[2x-z=0\] \[2z+y=-2\] So would we have to multiply the first equation so that the coefficient of x in the first equation is equal to the coefficient of x in the second equation?

OpenStudy (anonymous):

-2y-1z=-2

OpenStudy (anonymous):

I asked you a question. And that's incorrect.

OpenStudy (anonymous):

Okay. Forget about the third equation for now. Eliminate a variable in the first two equations.

OpenStudy (anonymous):

yes, you eliminate x by multiplying with -2

OpenStudy (anonymous):

YOu can just multiply by 2. That would make your life much easier. And then you can jsut subtract the two equations instead of adding them together....

OpenStudy (anonymous):

just*

OpenStudy (anonymous):

But if you want to do it that way, that's fine.

OpenStudy (anonymous):

Show me what you get when eliminate x.

OpenStudy (anonymous):

I'm still getting -2y-1z=-2 -2(x+y)=(1)-2 2x-z=0

OpenStudy (anonymous):

Multiply first equation by 2. \[2x+2y=2\] Subtract that with the second equation. \[(2x+2y)-(2x-z)=(2)-(0)\]

OpenStudy (anonymous):

What do you get now?

OpenStudy (anonymous):

2y-z=2

OpenStudy (anonymous):

GOod now eliminate one of the variables using the third equation and the equation you just got.

OpenStudy (anonymous):

Do exactly what I did. Just watch for your signs.

OpenStudy (anonymous):

y-z=0

OpenStudy (anonymous):

? You still have two variables.....I told you to eliminate one....

OpenStudy (anonymous):

sorry, messed up my variables

OpenStudy (anonymous):

2z+y=-2 2(-z+2y)=(2)2 y=\[\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

correct you aced it. Well done.

OpenStudy (anonymous):

you ca just sub that value into the equation to find z.

OpenStudy (anonymous):

so I got, \[(-\frac{ 3 }{ 5 },\frac{ 2 }{ 5 },-\frac{ 6 }{ 5 })\]

OpenStudy (anonymous):

It's 6/5

OpenStudy (anonymous):

Wait nvm you got it right.

OpenStudy (anonymous):

Well DOne. All correct.

OpenStudy (anonymous):

woo hoo

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

No worries.